# ARN: Analogical Reasoning on Narratives

### v3

@@ 4.1 Extracting Narrative Elements
A key benefit of the ePiC dataset is the already annotated proverbs associated with each narrative. [^2] We do not use span annotations in the ePiC dataset. Table 3 in the Appendix provides examples of narratives alongside their proverbs. We extract the other elements of a narrative using an LLM (GPT3.5; OpenAI, 2022) denoted by $llm(c,n)$ for extracting element $c$ from a narrative $n$. More concretely, we explicitly prompt the LLM asking it to extract the element, $c$ from the provided narrative, $n$, returning a Python List object. We opted to utilize GPT3.5 for this task to save both time and money, given its high accuracy in simple information extraction tasks (Wei et al., 2023b). Sample extracted elements from various narratives are shown in Table 1, which demonstrate the diversity of the subjects covered in the ePiC dataset. Across all narratives on average, we extracted 4.55 actions ($SD=1.90$), 2.44 goals ($SD=1.46$), 2.59 characters ($SD=0.96$), 1.71 relations ($SD=1.16$), and 3.90 locations ($SD=3.67$). After evaluating the extracted elements by a graduate student who was instructed to check whether the extracted elements were present in a subset of 100 random narratives, the precision of above $94\%$ was observed across all elements. We focus on precision in this pilot study since created mappings between two narratives based on a certain element, do not require identical elements. Additional details is provided in Section C.1.
@@ 4.4 Selecting & Generating Analogies
The narratives sharing the same proverb in the ePiC dataset are semantically far from each other by design. We, thus, leverage this property to establish far analogies between pairs of narratives that share the same proverb. To guarantee low surface similarity, for each query narrative $Q_{i}$ in $P$, we select the narrative with the same proverb that is semantically the farthest from $Q_{i}$ (using sentence BERT (Reimers and Gurevych, 2019) and cosine similarity) as a far analogy.
@@ 5 Experiments
We evaluate the analogical reasoning abilities of language models over narratives in $ARN$ using two settings. [^3] We did not find a difference between the models’ performance on all data points and data points where annotators had an agreement on. First, we report on their autonomous (Radford et al., 2019) analogical reasoning abilities to assess the emergent abilities of LLMs (Wei et al., 2022) and the latent analogical reasoning abilities they might have obtained during their training (see Section 5.1). Following a zero-shot formulation similar to Figure 1, we provide the model with the query narrative and the two candidate narratives, asking for the best analogical match to the query narrative. Asking for the most similar narrative yielded similar results (see Section D.3). We evaluate the autonomous analogical reasoning abilities of six LLMs: GPT3.5 (OpenAI, 2022), GPT4.0 (OpenAI, 2023), UnifiedQA (Khashabi et al., 2020), Llama-2 (Touvron et al., 2023), FlanT5 (Chung et al., 2022), and Macaw (Tafjord and Clark, 2021) using deterministic hyperparameters ($Temperature=0$). We also evaluate sentence BERT (SBERT, all-mpnet-base-v2; Reimers and Gurevych, 2019) by picking the narrative with a higher cosine similarity to the query narrative as a baseline. Additionally, in a more guided scenario, we experiment with a few-shot (in-context learning; Brown et al., 2020) setting to enhance models’ abilities with solved demonstrations (see Section 5.2). We evaluate Llama-2, GPT3.5, and GPT4.0 because of their in-context learning abilities and long context windows. Solved demonstrations are chosen randomly, and performance is reported as the average on three runs. For more details about the experimental setup and prompts, refer to Appendix B and Appendix C, respectively.
@@ 5.1 Autonomous Analogical Reasoning
Table 2: Accuracy (with a random baseline of $50\%$) of tested LLMs in zero-shot setting, on $ARN$ in four categories of far/near analogies facing far/near disanalogies, e.g., (near, far) indicates the setting where near analogies are alongside far disanalogies.
@@
In Table 2, we observe that models are not as good as humans at distinguishing analogies from distractors ($57$ vs $96\%$). On near analogies, the humans’ average performance is $97.9\%$, whereas the models’ performance averages $75.0\%$, which is halfway between random and human performance. Focusing merely on semantic similarity (using SBERT) already yields a good performance in this setting ($77.9\%$) which confirms the general similarity understanding of language models. Intuitively, the human-LLM gap is wider in solving far analogies where analogies have little surface similarity to query narratives. Although far analogies are slightly more difficult for humans, as evident in the small drop in performance, the drop is significant in LLMs. While humans’ average performance is $94.1\%$, models only achieve $39.6\%$ on average (worse than random performance) when solving far analogies. Similarly, we also observe that in both near and far analogies, as distractors get semantically more similar to the query narratives (far $\rightarrow$ near distractor), both LLMs’ and humans’ performance drop, with LLMs experiencing more notable drops. The most competitive model, GPT4.0, performs close to humans on near analogies; however, its performance is below random in far analogies. The relatively low performance of SOTA models underlines the challenging nature of analogical reasoning, even for the most capable LLMs.
@@
Comparing the performance of the models in different partitions of $ARN$, we observe the following patterns. All models perform well in choosing the near analogy over the far disanalogy ($76.9\%$), i.e., cases where analogies form surface mappings on top of system mappings and distractors do not form any mappings with the query narratives. When models are faced with two candidates with similar surface similarity (near-near or far-far), models’ performance drops significantly. Models score $73.2\%$ on average in a (near, near) setting which shows that models recognize system mappings in the presence of surface mappings, even though the distractor also forms surface mappings with the query narrative. Models perform much worse in the (far, far) setting (scoring only $45.3\%$), which shows their disability to recognize system mappings in the absence of surface mappings. This observation is further emphasized by the most challenging partition, where models need to prioritize far analogies over near distractors, which shows that models fail to prioritize system mappings over surface mappings (scoring $33.9\%$), regardless of their prompt instructions specifically asking for system mappings. Although solving examples in all partitions of the data requires analogical reasoning, the reasoning pattern specifically involved in each partition can be substituted by certain heuristics (e.g., the reasoner can achieve great scores by choosing the dissimilar narrative in the [far, near] condition without involving in analogical reasoning). We run our experiments completely isolated from each other, which makes the active development of such heuristics highly unlikely. However, it would be an interesting research direction to study the heuristics that models might use in analogical reasoning.
@@
As expected, the best-performing model is GPT4.0, with an average performance of $68.1\%$, and the worst-performing model is SBERT, with an average performance of $42.3\%$. GPT4’s performance is in line with prior findings about analogical reasoning being one of the emergent abilities of LLMs (Webb et al., 2023; Wei et al., 2022); however, its performance on $ARN$ is still halfway between random and human performance. Meanwhile, the poor performance of SBERT is expected, since it relies on surface semantic similarity between narratives, which makes it perform relatively well on recognizing near analogies but reaches only $1\%$ when prioritizing far analogies. While generally, we see that larger models perform better in the first three partitions, we observe that the best-performing model in the (far, near) setting is UnifiedQA-11B, even outperforming the GPT models with orders of magnitude more parameters. We attribute this curious observation to the smaller and more structured (instruction) fine-tuning that UnifiedQA has undergone and its alignment with the binary question-answering (QA) format of our task. To investigate this observation further, we also evaluate other QA models: FlanT5, Macaw, and smaller UnifiedQA variants. Aligned with our hypothesis, We observe that the other QA models also perform relatively well, scoring on par or better than the much larger foundational models: GPT3.5, GPT4.0, and Llama-2, on the (far, near) partition.
@@
We further study the distracting effect of each narrative element on the analogical reasoning abilities of the two best-performing models, GPT4.0 and UnifiedQA-11B. Figure 4 demonstrates the performance of models when distractors from each narrative element are excluded in the experiments, which we consider to be a proxy for the distraction effect of each narrative element — if excluding narrative elements $c_{1}$ from the distractor sampling leads to higher performance compared to excluding $c_{2}$, we infer that that $c_{1}$ has a higher distracting effect compared to $c_{2}$. For GPT4.0 (and all other models except UnifiedQA; see trends for other models in Section D.2), the aggregated distracting effect of all elements is stronger than individual elements, demonstrating the distinct effect of each element. Among the different elements, mappings based on actions lead to the most effective distractors both when detecting far and near analogies. In contrast, for UnifiedQA-11B, the aggregated effect of all surface mappings is not the most distracting, especially in far analogies where the model has the highest performance when all types of distractors are included. This observation, alongside the model having the highest performance in far analogies across all models, shows the model’s higher robustness for analogical reasoning, where it may ignore and even benefit from distracting surface mappings formed with disanalogies.
@@
Based on our observations on the free-text rationales and predictions by GPT4.0, the explanations generated by models are indeed based on high-level messages. However, these explanations can be incorrect in different ways that we iterate on some of the more salient ones:
@@
    The rationale is correct, while the prediction is incorrect. In the first example in Table 6 in the Appendix, the explanation is aligned with the correct choice, but the model is predicting the second narrative by mistake.
@@
    The provided explanation follows the main theme of the narratives and is aligned with the keywords of the query narrative and the chosen narrative. However, the explanation fails to focus on the high-level message of the query narrative. In the second example shown in Table 6 in the Appendix, the prediction is made based on narratives’ keywords. However, the high-level message that the correct answer is based on is not recognized by the model.
@@
    The rationale captures the high-level message of the query narrative and the chosen narrative but disregards the connection between the two narratives, i.e., the fact that a mapping should be formed between narratives. In the last example in Table 6 in the Appendix, the explanation and the gathered high-level message are aligned with the chosen narrative but do not hold true in the query narrative.
@@ 5.2 Guided Analogical Reasoning
A natural follow-up question is whether the performance of LLMs can be enhanced by guiding them through solved demonstrations, in comparison to a zero-shot setting. The performance of different LLMs in a few-shot setting is demonstrated in Figure 5 denoted by $\times$ signs, alongside their performance in a zero-shot setting where number of demonstrations is set to zero. When models are provided with solved demonstrations, their near-analogy skills degrade, whereas their far-analogy performance increases compared to the zero-shot setting. GPT4.0 is the most stable model, with its performance slightly increasing in far analogies (especially with more demonstrations), with a small drop in performance in near analogies. The other models, GPT3.5 and Llama-2, have more emphasized trends: they obtain a much larger gain on far analogies (close to 10 absolute points), while also losing much of their performance on near analogies, though part of the loss is mitigated by more demonstrations. Overall, these results suggest that the models do not efficiently leverage the solved demonstrations, and surprisingly, their performance in a few-shot setting is not necessarily better than zero-shot setting. We leave a deeper investigation of the models’ behavior to future work as they are beyond the scope of this paper.
@@
The models may be helped differentially by solved demonstrations that cover near or far analogies. To investigate this possibility, we test including either far or near analogies in the provided demonstrations to LLMs. Due to limitations in the expenses of using OpenAI’s API for GPT4.0, we conducted this experiment on GPT3.5 and Llama-2. The result of these experiments for GPT3.5 is presented in Figure 6 (Llama-2 showed similar trends). Overall, we observe similar patterns when using either near or far analogy demonstrations. Using either near or far analogy demonstrations improves the performance of models on detecting far analogies, while both demonstration types are detrimental to the models’ performance facing near analogies. However, increasing the number of far analogy demonstrations gradually improves the detection of near analogies, which suggests that far analogies can be slightly more helpful as demonstrations. The advantage of using far analogies as demonstrations can be explained by them being independent of the domains or specific content of analogs (Holyoak and Thagard, 1996).
@@
Solved demonstrations of analogies hold a promise to enhance performance while simultaneously being challenging to process, which may have led to the counter effects on model performance in Figure 5. Inspired by these observations, we investigate whether guiding the model to reason step-by-step can enable it to perform better. We employ chain-of-thought (COT) prompting (Wei et al., 2023a), given its positive impact on tasks that require complex reasoning. We experiment with Llama-2, GPT3.5, and GPT4.0 using $\{2,4,6\}$ demonstrations equally split between far and near analogies. In this step-by-step procedure, the model is instructed to extract the proverbial message of each narrative, followed by generating the final predictions. In general, COT demonstrations help the models more than normal demonstrations (see Figure 5). In near analogies, COT demonstrations still hurt the performance compared to the zero-shot setting, although this negative effect is less strong than in the normal few-shot setting. However, in far analogies, COT demonstrations are more beneficial than both the normal few-shot demonstrations and no demonstrations (zero-shot setting). Here, GPT-based models gain up to 30 absolute points compared to the zero-shot setting. The models’ performance not improving for near analogies is likely because the models are already using surface similarity cues to solve the task, and COT’s reasoning chains are not recognized as critical by the models. The improvements in far analogies are likely because, in the absence of surface similarity, models need to look for elaborate mappings, and COT can facilitate this, by breaking the task into smaller, more manageable parts.

### v4

@@ 4.1 Extracting Narrative Elements
A key benefit of the ePiC dataset is the already annotated proverbs associated with each narrative. [^2] We do not use span annotations in the ePiC dataset. Table 3 in the Appendix provides examples of narratives alongside their proverbs. We extract the other elements of a narrative using an LLM (GPT3.5; OpenAI, 2022) denoted by $llm(c,n)$ for extracting element $c$ from a narrative $n$. More concretely, we explicitly prompt the LLM, asking it to extract the element, $c$, from the provided narrative, $n$, returning a Python List object. We opted to utilize GPT3.5 for this task to save both time and money, given its high accuracy in simple information extraction tasks (Wei et al., 2023). Sample extracted elements from various narratives are shown in Table 1, which demonstrate the diversity of the subjects covered in the ePiC dataset. Across all narratives, on average, we extracted 4.55 actions ($SD=1.90$), 2.44 goals ($SD=1.46$), 2.59 characters ($SD=0.96$), 1.71 relations ($SD=1.16$), and 3.90 locations ($SD=3.67$). After evaluating the extracted elements by a graduate student who was instructed to check whether the extracted elements were present in a subset of 100 random narratives, the precision of above $94\%$ was observed across all elements. We focus on precision in this pilot study since created mappings between two narratives based on a certain element do not require identical elements. Additional details are provided in Section C.1.
@@ 4.4 Selecting & Generating Analogies
The narratives sharing the same proverb in the ePiC dataset are semantically far from each other by design. We, thus, leverage this property to establish far analogies between pairs of narratives that share the same proverb. To guarantee low surface similarity, for each query narrative $Q_{i}$ in $P$, we select the narrative with the same proverb that is semantically the farthest from $Q_{i}$ (using sentence BERT; Reimers and Gurevych, 2019; and cosine similarity) as a far analogy.
@@ 5 Experiments
We evaluate the analogical reasoning abilities of language models over narratives in $ARN$ using two settings. [^3] We did not find a difference between the models’ performance on all data points and data points where annotators had an agreement on. First, we report on their autonomous (Radford et al., 2019) analogical reasoning abilities to assess the emergent abilities of LLMs (Wei et al., 2022a) and the latent analogical reasoning abilities they might have obtained during their training (see Section 5.1). Following a zero-shot formulation similar to Figure 1, we provide the model with the query narrative and the two candidate narratives, asking for the best analogical match to the query narrative. Asking for the most similar narrative yielded similar results (see Section D.3). We evaluate the autonomous analogical reasoning abilities of six LLMs: GPT3.5 (OpenAI, 2022), GPT4.0 (OpenAI et al., 2024), UnifiedQA (Khashabi et al., 2020), Llama-2 (Touvron et al., 2023), FlanT5 (Chung et al., 2024), and Macaw (Tafjord and Clark, 2021) using deterministic hyperparameters ($Temperature=0$). We also evaluate sentence BERT (SBERT, all-mpnet-base-v2; Reimers and Gurevych, 2019) by picking the narrative with a higher cosine similarity to the query narrative as a baseline. Additionally, in a more guided scenario, we experiment with a few-shot (in-context learning; Brown et al., 2020) setting to enhance models’ abilities with solved demonstrations (see Section 5.2). We evaluate Llama-2, GPT3.5, and GPT4.0 because of their in-context learning abilities and long context windows. Solved demonstrations are chosen randomly, and performance is reported as the average on three runs. For more details about the experimental setup and prompts, refer to Appendix B and Appendix C, respectively.
@@ 5.1 Autonomous Analogical Reasoning
Table 2: Accuracy (with a random baseline of $50\%$) of tested LLMs in zero-shot setting, on $ARN$ in four categories of far/near analogies facing far/near disanalogies, e.g., (near, far) indicates the setting where near analogies are alongside far disanalogies. Boldfaced numbers in each partition show the best-performing model in that partition.
@@
In Table 2, we observed that models are not as good as humans at distinguishing analogies from distractors ($57$ vs $96\%$). On near analogies, the humans’ average performance was $97.9\%$, whereas the models’ performance averaged $75.0\%$, which is halfway between random and human performance. Focusing merely on semantic similarity (using SBERT) already yielded a good performance in this setting ($77.9\%$) which confirms the general similarity understanding of language models. Intuitively, the human-LLM gap was wider in solving far analogies where analogies have little surface similarity to query narratives. Although far analogies were slightly more difficult for humans, as evident in the small drop in performance, the drop was significant in LLMs. While humans’ average performance was $94.1\%$, models only achieved $39.6\%$ on average (worse than random performance) when solving far analogies. Similarly, we also observed that in both near and far analogies, as distractors get semantically more similar to the query narratives (far $\rightarrow$ near distractor), both LLMs’ and humans’ performance drop, with LLMs experiencing more notable drops. The most competitive model, GPT4.0, performed close to humans on near analogies; however, its performance was below random in far analogies. The relatively low performance of SOTA models underlines the challenging nature of analogical reasoning, even for the most capable LLMs.
@@
Comparing the performance of the models in different partitions of $ARN$, we observed the following patterns. All models performed well in choosing the near analogy over the far disanalogy ($76.9\%$), i.e., cases where analogies form surface mappings on top of system mappings and distractors do not form any mappings with the query narratives. Facing two candidates with similar surface similarity (near-near or far-far), models’ performance dropped significantly. Models scored $73.2\%$ on average in a (near, near) setting which shows that models recognize system mappings in the presence of surface mappings, even though the distractor also forms surface mappings with the query narrative. Models performed much worse in the (far, far) setting (scoring only $45.3\%$), which shows their disability to recognize system mappings in the absence of surface mappings. This observation was further emphasized by the most challenging partition, where models needed to prioritize far analogies over near distractors, as models failed to prioritize system mappings over surface mappings (scoring $33.9\%$), regardless of their prompt instructions specifically asking for system mappings. Note that although solving examples in all partitions of the data requires analogical reasoning, the reasoning pattern specifically involved in each partition can be substituted by certain heuristics (e.g., the reasoner can achieve great scores by choosing the dissimilar narrative in the [far, near] condition without involving analogical reasoning). We run our experiments completely isolated from each other, which makes the active development of such heuristics highly unlikely. However, it would be an interesting research direction to study the heuristics that models might use in analogical reasoning.
@@
As expected, the best-performing model was GPT4.0, with an average performance of $68.1\%$, and the worst-performing model was SBERT, with an average performance of $42.3\%$. GPT4’s performance was in line with prior findings about analogical reasoning being one of the emergent abilities of LLMs (Webb et al., 2023; Wei et al., 2022a); however, its performance on $ARN$ was still halfway between random and human performance. Meanwhile, the poor performance of SBERT was expected since it relies on surface semantic similarity between narratives, which makes it perform relatively well on recognizing near analogies but reaches only $1\%$ when prioritizing far analogies. While generally, we saw that larger models perform better in the first three partitions, we observed that the best-performing model in the (far, near) setting is UnifiedQA-11B, even outperforming the GPT models with orders of magnitude more parameters. We attribute this curious observation to the smaller and more structured (instruction) fine-tuning that UnifiedQA has undergone and its alignment with the binary question-answering (QA) format of our task. To investigate this observation further, we also evaluated other QA models: FlanT5, Macaw, and smaller UnifiedQA variants. Aligned with our hypothesis, We observed that the other QA models also perform relatively well, scoring on par or better than the much larger foundational models: GPT3.5, GPT4.0, and Llama-2, on the (far, near) partition.
@@
We further studied the distracting effect of each narrative element on the analogical reasoning abilities of the two best-performing models, GPT4.0 and UnifiedQA-11B. Figure 4 demonstrates the performance of models when distractors from each narrative element were excluded in the experiments, which we consider to be a proxy for the distraction effect of each narrative element — if excluding narrative elements $c_{1}$ from the distractor sampling leads to higher performance compared to excluding $c_{2}$, we infer that that $c_{1}$ has a higher distracting effect compared to $c_{2}$. For GPT4.0 (and all other models except UnifiedQA; see trends for other models in Section D.2), the aggregated distracting effect of all elements was stronger than individual elements, demonstrating the distinct effect of each element. Among the different elements, mappings based on actions led to the most effective distractors both when detecting far and near analogies. In contrast, for UnifiedQA-11B, the aggregated effect of all surface mappings was not the most distracting, especially in far analogies where the model had the highest performance when all types of distractors were included. This observation, alongside the model having the highest performance in far analogies across all models, shows the model’s higher robustness for analogical reasoning, where it may ignore and even benefit from distracting surface mappings formed with disanalogies.
@@
Based on our observations on the free-text rationales and predictions by GPT4.0, the explanations generated by models were indeed based on high-level messages. However, these explanations could be incorrect in different ways that we iterate on some of the more salient ones:
@@
    The rationale was correct, while the prediction was incorrect. In the first example in Table 6 in the Appendix, the explanation was aligned with the correct choice, but the model predicted the second narrative by mistake.
@@
    The provided explanation followed the main theme of the narratives and was aligned with the keywords of the query narrative and the chosen narrative. However, the explanation failed to focus on the high-level message of the query narrative. In the second example shown in Table 6 in the Appendix, the prediction was made based on narratives’ keywords. However, the high-level message that the correct answer is based on was not recognized by the model.
@@
    The rationale captured the high-level message of the query narrative and the chosen narrative but disregarded the connection between the two narratives, i.e., the fact that a mapping should be formed between narratives. In the last example in Table 6 in the Appendix, the explanation and the gathered high-level message were aligned with the chosen narrative but did not hold true in the query narrative.
@@ 5.2 Guided Analogical Reasoning
A natural follow-up question is whether the performance of LLMs can be enhanced by guiding them through solved demonstrations in comparison to a zero-shot setting. The performance of different LLMs in a few-shot setting is demonstrated in Figure 5 denoted by $\times$ signs, alongside their performance in a zero-shot setting where the number of demonstrations is set to zero. When models were provided with solved demonstrations, their near-analogy skills degraded, whereas their far-analogy performance increased compared to the zero-shot setting. GPT4.0 was the most stable model, with its performance slightly increasing in far analogies (especially with more demonstrations), with a small drop in performance in near analogies. The other models, GPT3.5 and Llama-2, had more emphasized trends: they obtained a much larger gain on far analogies (close to 10 absolute points) while also losing much of their performance on near analogies, though more demonstrations mitigated part of the loss. Overall, these results suggest that the models do not efficiently leverage the solved demonstrations, and surprisingly, their performance in a few-shot setting is not necessarily better than zero-shot setting. We leave a deeper investigation of the models’ behavior to future work as they are beyond the scope of this paper.
@@
The models may use solved demonstrations that cover near or far analogies differentially, which prompted us to test including either far or near analogies in the provided demonstrations to LLMs. Due to limitations in the expenses of using OpenAI’s API for GPT4.0, we conducted this experiment on GPT3.5 and Llama-2. The result of these experiments for GPT3.5 is presented in Figure 6 (Llama-2 showed similar trends). Overall, we observed similar patterns when using either near or far analogy demonstrations. Using either near or far analogy demonstrations improved the performance of models on detecting far analogies, while both demonstration types were detrimental to the models’ performance facing near analogies. However, increasing the number of far analogy demonstrations gradually improved the detection of near analogies, which suggests that far analogies can be slightly more helpful as demonstrations. The advantage of using far analogies as demonstrations can be explained by them being independent of the domains or specific content of analogs (Holyoak and Thagard, 1996).
@@
Solved demonstrations of analogies hold a promise to enhance performance while simultaneously being challenging to process, which may have led to the counter effects on model performance in Figure 5. Inspired by these observations, we investigated whether guiding the model to reason step-by-step can enable it to perform better. We employed Chain-of-Thought (CoT) prompting (Wei et al., 2022b), given its positive impact on tasks that require complex reasoning. We experimented with Llama-2, GPT3.5, and GPT4.0 using $\{2,4,6\}$ demonstrations equally split between far and near analogies. In this step-by-step procedure, the model was instructed to extract the proverbial message of each narrative, followed by generating the final predictions. In general, CoT demonstrations helped the models more than normal demonstrations (see Figure 5). In near analogies, CoT demonstrations still hurt the performance compared to the zero-shot setting, although this negative effect was less strong than in the normal few-shot setting. However, in far analogies, CoT demonstrations were more beneficial than both the normal few-shot demonstrations and no demonstrations (zero-shot setting). Here, GPT-based models gained up to 30 absolute points compared to the zero-shot setting. The models’ performance not improving for near analogies is likely because the models are already using surface similarity cues to solve the task, and CoT’s reasoning chains are not recognized as critical by the models. The improvements in far analogies are likely because models need to look for elaborate mappings without surface similarity, and CoT can facilitate this by breaking the task into smaller, more manageable parts.